OpenStudy (anonymous):

classify each system as a dependent, independent, or inconsistent. {-2x-y=9} {3x-4y=-8} both are in same bracket.

OpenStudy (anonymous):

Multiply the top equation by \(-4\) and add them.

OpenStudy (anonymous):

\[ {-2x-y=9 \brace 3x-4y=-8 } \]

OpenStudy (anonymous):

\[ { 8x+4y=-36 \brace 3x-4y=-8 } \]

OpenStudy (anonymous):

Add them an they will cancel.

OpenStudy (anonymous):

so it will be a inconsistent ! correct

OpenStudy (anonymous):

Well, it will be independent.

OpenStudy (anonymous):

inconsistent means there is no solution.

OpenStudy (usukidoll):

inconsistent system has the following format...  where * is any number